OpenStudy (anonymous):

Jessica skated 4 laps in 82 seconds. Which of the following is an equivalent rate of skating

OpenStudy (sleepyjess):

Welcome to OpenStudy! It looks like there should be some options. Could you post them? :)

OpenStudy (texaschic101):

4/82....basically you are looking for equivalent fractions. 4/82 = 2/41 ...so it is the same as swimming 2 laps in 41 seconds or 4/82 * 2/2 = 8/164...same as swimming 8 laps in 164 seconds or (2 minutes & 44 seconds) or 4/82 * 3/3 = 12/246...same as swimming 12 laps in 246 seconds or (4 min & 6 seconds) I could go on for a while...it would be easier if you posted the answer choices.